# FakeShield: Explainable Image Forgery Detection and Localization via Multi-modal Large Language Models

### v3

@@ A.4 Prompts
A.5 Examples
Comparison of subjective results of mainstream M-LLM: As mentioned in Section 4.3, we selected some M-LLM output samples, as shown in Figures 9 and 10.
FakeShield output subjective samples: We selected several results from FakeShield’s testing on PhotoShop, DeepFake, and AIGC-Editing datasets, as displayed in Figures 1112131415 and 16.
MMTD-Set data set example: We select some samples from the MMTD-Set data set and display them in Figures 1718, and 19.
@@ A.5 Examples
Figure 11: Dialogue examples of FakeShield when receiving PhotoShop tampered images. The first image comes from the CASIA1+ (Dong et al., 2013) dataset and has been tampered with by Copy-Move. The yellow petals in the center of the picture are the parts pasted later. The second picture comes from the CASIA1+ (Dong et al., 2013) dataset and has been tampered with by Splicing. The penguin in the center of the picture is the part pasted later.
@@
Figure 14: Dialogue examples of FakeShield when receiving DeepFake tampered images. The three images all come from the DFFD (Dang et al., 2020) dataset and have been tampered with by FaceAPP (FaceApp Limited, 2017). The first picture has a beard added to the chin, the first picture has glasses added, and the third picture has some hair added to the right side of the face.

### v4

@@ A.6 Prompts
A.7 Examples
Comparison of subjective results of mainstream M-LLM: As mentioned in Section 4.3, we selected some M-LLM output samples, as shown in Figures 10 and 11.
FakeShield output subjective samples: We selected several results from FakeShield’s testing on PhotoShop, DeepFake, and AIGC-Editing datasets, as displayed in Figures 1213141516 and 17.
MMTD-Set data set example: We select some samples from the MMTD-Set data set and display them in Figures 1819, and 20.
@@ A.7 Examples
Figure 12: Dialogue examples of FakeShield when receiving PhotoShop tampered images. The first image comes from the CASIA1+ (Dong et al., 2013) dataset and has been tampered with by Copy-Move. The yellow petals in the center of the picture are the parts pasted later. The second
picture comes from the CASIA1+ (Dong et al., 2013) dataset and has been tampered with by Splicing. The penguin in the center of the picture is the part pasted later.
@@
Figure 15: Dialogue examples of FakeShield when receiving DeepFake tampered images. The three images all come from the DFFD (Dang et al., 2020) dataset and have been tampered with by FaceAPP (FaceApp Limited, 2017). The first picture has a beard added to the chin, the second picture has glasses added, and the third picture has some hair added to the right side of the face.